OpenStudy (kamille):

There 14 lottery tickets and 6 of them are lucky(you can win something). How many ways are there to 2 lucky tickets, if you can take 5 of 14. (the answer is 85). Any tips?

OpenStudy (kamille):

to take 2 lucky tickets*

OpenStudy (kamille):

@Mertsj

OpenStudy (kamille):

i can try to rewrite a problem, if you dont understand what i need to get

OpenStudy (anonymous):

Try, then :)

OpenStudy (kamille):

There is 14 lottery tickets and 6 of them are lucky (you can win something). Dou you understand this?

OpenStudy (anonymous):

Yes...

OpenStudy (kamille):

so, you pick 5 of 14 tickets. Understandable?

OpenStudy (anonymous):

Yes.

OpenStudy (kamille):

How many ways are to pick that 5 tickets, if there must be 2 lucky ones?

OpenStudy (kamille):

the answer is 85, if it helps.

OpenStudy (anonymous):

At least 2, or only 2?

OpenStudy (kamille):

only

OpenStudy (kamille):

oh, and i gave wrong answer, it is 840

OpenStudy (anonymous):

:/

OpenStudy (kamille):

://

OpenStudy (anonymous):

Well, It was hard earlier, because you said the answer was 85... which seemed too small.

OpenStudy (kamille):

so,is it easier now?

OpenStudy (anonymous):

Much.

OpenStudy (anonymous):

It's basically the product of how many ways you can pick two lucky tickets and how many ways you can pick three not-lucky tickets :)

OpenStudy (kamille):

Can you explain me step by step? I still dont get it :/

OpenStudy (anonymous):

Well, you are going to pick five tickets, and you want to know how many ways you can pick a combination of five where exactly two are lucky, right? Well, that means exactly three are not-lucky.

OpenStudy (anonymous):

\[C^6_2\times C^{14-6}_3=\frac{6\times5\times\cancel{4!}}{\cancel{4!}\times2}\times\frac{8\times7\times6\times\cancel{5!}}{\cancel{5!}\times3\times2}=840\]

OpenStudy (anonymous):

So, it boils down to how many ways you can pick two lucky tickets, and there are 6 of them. So it'd be 6C2

OpenStudy (anonymous):

times... the number of ways you can pick three not-lucky tickets, and there are 8 of them. So it'd be 8C3

OpenStudy (kamille):

oh, okay! Thank you a lot! I have another problem, with "at least", what is different about it?

OpenStudy (anonymous):

We'll see :P